# Mitigating Hallucinations in LM-Based TTS Models via Distribution Alignment Using GFlowNets

### v2

@@ Abstract
Language Model (LM)-based Text-to-Speech (TTS) systems often generate hallucinated speech that deviates from input text. Existing mitigation strategies either demand excessive training resources or introduce significant inference latency. In this paper, we propose GFlOwNet-guided distribution AlignmenT (GOAT) for LM-based TTS, a post-training framework that mitigates hallucinations without relying on massive resources or inference cost. Specifically, we first conduct an uncertainty analysis, revealing a strong positive correlation between hallucination and model uncertainty. Based on this, we reformulate TTS generation as a trajectory flow optimization problem and introduce an enhanced Subtrajectory Balance objective together with a sharpened internal reward as target distribution. We further integrate reward temperature decay and learning rate optimization for stability and performance balance. Extensive experiments show that GOAT reduce over 50% character error rates on challenging test cases and lowering uncertainty by up to 58%, demonstrating its strong generalization ability and effectiveness.
@@ 1 Introduction
Nevertheless, several studies Ji et al. (2023); Chuang et al. (2023) in the field of natural language processing have shown that language models are prone to hallucination, particularly when predicting tokens that convey factual information. As shown in Figure 1. This phenomenon has unfortunately been inherited by LM-based TTS models, manifesting as generated speech that may deviate from the ground-truth text. Such issues become more pronounced when generating long or complex sentences, with errors such as mispronunciations, word omissions, and unnatural repetitions occurring more frequently.
@@ 2.1 Hallucination Detection
Recent studies Huang et al. (2024a); Ma et al. (2025) have analyzed the hallucination problem of language models in text-based tasks from the perspective of model uncertainty. Inspired by this, we adopt the entropy-based metric for speech hallucination detection. More detailed discussion can be found in Appendix A.1. Given that speech tokens have significantly lower information density compared to text tokens, multiple tokens must be aggregated to represent the pronunciation of a single word. With this in mind, we assess the uncertainty of LM-based TTS models across three dimensions: token level, word level, and utterance level.
@@
where $|C|$ denotes the vocabulary size. We extract the left and right boundaries $i$ and $j$ of each word $W$ within the token sequence to compute the word-level uncertainty $H_{\text{word}}(W_{i:j})$, while the utterance-level uncertainty $H_{\text{utterance}}(S)$ is defined as the average uncertainty over all tokens. They can be formally defined as follows:
$H_{\text{word}}(W_{i:j})=\frac{1}{j-i}\sum_{k=i}^{j}H_{\text{token}}(P_{k})$ (2)
@@
where $|S|$ denotes the sequence length. More detailed discussion can be found in Appendix A.2. Intuitively quantifying model uncertainty facilitates the exploration of its potential connection to hallucination.
@@ 2.2 Empirical Analysis
We conducted experiments on SeedTTS-Eval benchmark Anastassiou et al. (2024) utilizing CosyVoice2 Du et al. (2024b). To reveal potential hallucinations, we select 200 samples from the test-hard subset and generate speech using stochastic multinomial sampling. Paraformer-zh Gao et al. (2022) is employed to perform automatic speech recognition (ASR) to compute the word error rate (WER), which serves as the hallucination proxy.
As shown in Figure 2, we annotate the utterance-level uncertainty and the corresponding WER for all samples, and perform linear regression. We observe that utterances with low WER typically exhibit lower uncertainty, and vice versa. To enhance the credibility of the results, we further compute the Pearson correlation coefficient and the Spearman rank correlation coefficient, which are 0.636 and 0.649 respectively ($p<1\mathrm{E}-4$), revealing statistically significant positive correlations. The detailed analysis of token-level uncertainty and word-level uncertainty is provided in Appendix A.3.
Figure 2: The relationship between utterance-level uncertainty and Log-WER. The red line represents the linear regression fit.
@@ 3.2.2 Internal Reward
$R(\mathbf{a}_{k}|\mathbf{q})=p_{LM}(\mathbf{a}_{k}|\mathbf{q})$ (13)
@@ 4.1 Datasets & Baseline
Datasets We randomly select 1000 samples from LibriTTS Zen et al. (2019) and WenetSpeech4TTS Ma et al. (2024) for training on Chinese and English. For evaluation, we use the SeedTTS-Eval benchmark Anastassiou et al. (2024), with further details provided in Appendix C..
@@ 4.3 Sampling Methods
The sampling strategy employed in CosyVoice2 is Repetition Aware Sampling (RAS), a combination of top-p and top-k. For a more fine-grained comparison, we introduce random multinomial sampling (RMS). Additionally, to verify the difference discussed in Section 3.2.2, we also introduce baseline sampling with the same low temperature as the reward function for comparison.
@@ 4.4.1 Metrics
Table 1: Evaluation results across models. CV2-GOAT-zh/en/mix: Models fine-tuned on Chinese/English/Mix of two datasets; 1500S/2500S denotes total steps in Learning Rate Optimization. SM: Sampling Method; LT prefix: sampling method using Low-Temperature.
@@ 4.5 In-depth Analysis
Figure 5: Comparison of the utterance-level uncertainty and log WER between the two models. For those who have zero WER, we set the Log WER to 0.
@@
Uncertainty Analysis We also evaluate the effectiveness of GOAT from the perspective of uncertainty. Leveraging the utterance-level uncertainty defined in Formula 3, we compare CV2-GOAT-zh(2500S) with the baseline as shown in Figure 2. The results in Figure 5 reveal that the aligned model exhibits significantly lower utterance-level uncertainty than the baseline. Additionally, the overall data cluster shifts leftward, indicating reduced WER (as confirmed in Section 4.4.2), demonstrating effective hallucination suppression.
@@ A.2 Hallucination Detection Methods
Figure 6: Model uncertainty quantification methods on token, word and utterance levels by entropy.
@@
Equation (16) represents the uncertainty analysis at the token level. However, in speech token sequences, multiple tokens are often required to represent the pronunciation of a single word. For a word $W_{ij}$ composed of the token sequence $a_{i},a_{i+1},\dots,a_{j}$, the uncertainty $H_{\text{word}}(W_{ij})$ is computed as the average uncertainty of the tokens constituting the word. Similarly, for an utterance $S$ with a token sequence length of $|S|$, the uncertainty $H_{\text{utterance}}(S)$ is calculated as the mean uncertainty of all tokens in the utterance. The formulas for word-level and utterance-level uncertainty are as follows:
$H_{\text{word}}(W_{ij})=\frac{1}{j-i}\sum_{k=i}^{j}H_{\text{token}}(P_{a_{k}})$ (17)
@@
An illustration of our detection method is provided in Figure 6. Based on In the context of LLM generation tasks, entropy can be interpreted as a measure of the model’s uncertainty regarding its output. When the model is highly confident in predicting a specific token, the resulting probability distribution exhibits low entropy, indicating a stable generation process. Conversely, a high-entropy distribution suggests that the model is uncertain among multiple candidate tokens, leading to increased sampling variability and potential susceptibility to randomness-induced fluctuations.
@@ A.3 Detailed Fine-grained Hallucination Phenomena Analysis with Uncertainty
For fine-grained analysis, we conduct token-level uncertainty during utterance-level uncertainty computation, and visualized token-wise uncertainty via line plots for each test utterance to conduct fine-grained analysis. By leveraging temporal alignment from ASR outputs, we identified token subsequences corresponding to individual words. For word-level uncertainty, we computed the average uncertainty of constituent tokens and generated word-wise uncertainty line plots.
@@
Figure 7: One classic case of Mispronounced/Incorrect Words, with red dashed boxes hightling error regions.
Mispronounced/Incorrect Words In these cases, the model generates speech containing lexically incorrect terms or phonetic errors. Since each word in speech token sequences is represented by multiple tokens, such errors manifest as faulty subsequence generation, indicating hallucinatory behavior. As is visualized in Figure 7, both token-level and word-level uncertainties are significantly elevated in these cases.
@@
Repetition/Deletion Errors These kind of errors occur when the model fails to handle repetitive text, resulting in missing or extra word repetitions. As is visualized in Figure 8, some errors arise from high uncertainty (e.g., skipped words in complex repetitions), while others stem from overconfidence—where uncertainty decreases despite errors.
Figure 9: Two classic cases of Mid-Generation Collapse, with red dashed boxes hightling error regions. (a) A mid-generation collapse occurred in the speech token sequence (b) Starting from a certain step in the middle, the generated content becomes entirely incoherent until termination.
@@
Figure 10: One classic case of Semantic Divergence, with red dashed boxes hightling error regions. The first error region includes word mispronunciations, and the second large region is inconsistent with target text although preserving semantically related contents.
Semantic Divergence This involves semantically plausible but textually mismatched content, typically in long sentences.As is visualized in Figure 10, while the generated speech remains intelligible and semantically related to targets, textual alignment breaks down, which represents the occurrence of hallucinations. Relatively large uncertainty is observed despite the superficial coherence of outputs.
Figure 11: Three classic cases of Prosody errors, with red dashed boxes hightling error regions. (a) The error region contains two characters with a gap that should be positioned between them. (b) In this error region, the separated characters should be coherently pronounced, followed by a brief pause. (c) Both larger gaps are redundant silences and should not occur in this error region.
Prosody Errors Prosody errors include misplaced pauses or redundant silences that may impair content comprehension. As is visualized in Figure 11, unlike other categories, no consistent uncertainty pattern emerges for prosodic hallucinations. Only part of high uncertainty tokens are observed in some pauses or words nearby. This complexity suggests the need for specialized uncertainty metrics tailored to prosodic representation learning.
@@ Appendix C Datasets
Evaluation Dataset Following the methodology of CosyVoice 2, we used the official seed-tts-eval test setAnastassiou et al. (2024), which contains three sub datasets, including approximately 1,000 English test samples (test-en) from the Common Voice dataset, and 2,000 Chinese test samples (test-zh) from the DiDiSpeech-2 dataset. Additionally, there are around 400 challenging samples in the case test-hard subset, which includes difficult synthesis targets such as text repetition, tongue twisters, phonetically similar texts, and commonly mispronounced words. We use the same testing methodology and dataset as CosyVoice 2 to demonstrate the improvements made by our approach in enhancing model performance.
@@ Appendix D Implementation Details
In this experiment, we provide two sets of hyperparameter configurations, which differ primarily in the total number of steps in the Learning Rate Optimization (LRO) process. Specifically, for all experiments, the total training rounds are fixed at 15, which corresponds to approximately 3500 global steps. As described in Section  3.3, to balance model performance and training stability while suppressing reward hacking, we set the minimum reward temperature to 0.825 and the maximum learning rate to 1e-5. Two configurations with slight differences in LRO are provided, where the warm-up phase is fixed at 20 steps. We adjust the remaining cosine annealing steps to achieve total LRO steps of 1500 and 2500, corresponding to more stable training and stronger model performance respectively.
@@ Appendix G Inference Latency Analysis
We use Real-Time Factor (RTF) for inference latency analysisis. RTF is a widely used metric in the speech generation field to measure the efficiency of a model in terms of its processing speed. It is defined as the ratio between the time taken for the model to generate speech and the duration of the generated audio. Specifically, the RTF is calculated as:

### v3

@@ Abstract
Language Model (LM)-based Text-to-Speech (TTS) systems often generate hallucinated speech that deviates from input text. Existing mitigation strategies either demand excessive training resources or introduce significant inference latency. In this paper, we propose GFlOwNet-guided distribution AlignmenT (GOAT) for LM-based TTS, a post-training framework that mitigates hallucinations without relying on massive resources or inference cost. Specifically, we first conduct an uncertainty analysis, revealing a strong positive correlation between hallucination and model uncertainty. Based on this, we reformulate TTS generation as a trajectory flow optimization problem and introduce an enhanced Subtrajectory Balance objective together with a sharpened internal reward as target distribution. We further integrate reward temperature decay and learning rate optimization for stability and performance balance. Extensive experiments show that GOAT reduce over 50% character error rates on challenging test cases and lowering uncertainty by up to 58%, demonstrating its strong generalization ability and effectiveness. Code: https://github.com/lotuscarvedlife/GOAT
@@ 1 Introduction
Nevertheless, several studies Ji et al. (2023); Chuang et al. (2023) in the field of natural language processing have shown that language models are prone to hallucination, particularly when predicting tokens that convey factual information, as shown in Figure 1. This phenomenon has unfortunately been inherited by LM-based TTS models, manifesting as generated speech that may deviate from the ground-truth text. Such issues become more pronounced when generating long or complex sentences, with errors such as mispronunciations, word omissions, and unnatural repetitions occurring more frequently.
@@ 2.1 Hallucination Detection
Recent studies Huang et al. (2024a); Ma et al. (2025) have analyzed the hallucination problem of language models in text-based tasks from the perspective of model uncertainty. Inspired by this, we adopt the entropy-based metric for speech hallucination detection. More detailed discussion can be found in Appendix A.1. Given that speech tokens have significantly lower information density compared to text tokens, multiple tokens must be aggregated to represent the pronunciation of a single character. With this in mind, we assess the uncertainty of LM-based TTS models across three dimensions: token level, character level, and utterance level.
@@
where $|C|$ denotes the vocabulary size. We extract the left and right boundaries $i$ and $j$ of each character $W$ within the token sequence to compute the character-level uncertainty $H_{\text{character}}(W_{i:j})$, while the utterance-level uncertainty $H_{\text{utterance}}(S)$ is defined as the average uncertainty over all tokens. They can be formally defined as follows:
$H_{\text{character}}(W_{i:j})$ $=\frac{1}{j-i}\sum_{k=i}^{j}H_{\text{token}}(P_{k})$ (2)
@@
where $|S|$ denotes the sequence length. More discussion can be found in Appendix A.2. Intuitively quantifying model uncertainty facilitates the exploration of its potential connection to hallucination.
@@ 2.2 Empirical Analysis
We conducted experiments on SeedTTS-Eval benchmark Anastassiou et al. (2024) utilizing CosyVoice 2 Du et al. (2024b). To reveal potential hallucinations, we select 200 samples from the test-hard subset and generate speech using stochastic multinomial sampling. Paraformer-zh Gao et al. (2022) is employed to perform automatic speech recognition (ASR) to compute the character error rate (CER), which serves as the hallucination proxy.
As shown in Figure 2, we annotate the utterance-level uncertainty and the corresponding CER for all samples, and perform linear regression. We observe that utterances with low CER typically exhibit lower uncertainty, and vice versa. To enhance the credibility of the results, we further compute the Pearson correlation coefficient and the Spearman rank correlation coefficient, which are 0.636 and 0.649 respectively ($p<1\mathrm{E}-4$), revealing statistically significant positive correlations. More discussion and detailed analysis of token-level/character-level uncertainty is provided in Appendix A.3.
Figure 2: The relationship between utterance-level uncertainty and log CER. The red line represents the linear regression fit.
@@ 3.2.2 Internal Reward
$R(\mathbf{a}_{k}|\mathbf{q})=p_{\text{LM}}(\mathbf{a}_{k}|\mathbf{q})$ (13)
@@ 4.1 Datasets & Baseline
Datasets We randomly select 1000 samples from LibriTTS Zen et al. (2019) and WenetSpeech4TTS Ma et al. (2024) for training on Chinese and English. For evaluation, we use the SeedTTS-Eval benchmark Anastassiou et al. (2024), with further details provided in Appendix C.
@@ 4.3 Sampling Methods
We adopt the Repetition Aware Sampling (RAS) strategy from CosyVoice 2 under its default hyperparameters (see Appendix D for details). For a more fine-grained comparison, we introduce random multinomial sampling (RMS). Additionally, to verify the difference discussed in Section 3.2.2, we also include a baseline that applies the same low temperature as used in the reward function.
@@ 4.4.1 Metrics
Table 1: Evaluation results across models. CV2-GOAT-zh/en/mix: models fine-tuned on Chinese/English/Mix of two datasets; 1500S/2500S : total steps in Learning Rate Optimization. SM: Sampling Method; LT prefix: sampling method using Low-Temperature.
@@ 4.5 In-depth Analysis
Figure 5: Comparison of the utterance-level uncertainty and log CER between the two models. For those who have zero CER, we set the log CER to 0.
@@
Uncertainty Analysis We also evaluate the effectiveness of GOAT from the perspective of uncertainty. Leveraging the utterance-level uncertainty defined in Formula 3, we compare CV2-GOAT-zh(2500S) with the baseline as shown in Figure 2. The results in Figure 5 reveal that the aligned model exhibits significantly lower utterance-level uncertainty than the baseline. Additionally, the overall data cluster shifts leftward, indicating reduced CER (as confirmed in Section 4.4.2), demonstrating effective hallucination suppression.
@@ 7 Limitations
8 Acknowledgments
This research is supported by the National Natural Science Foundation of China under Grant No. 62376071.
@@ A.2 Hallucination Detection Methods
Figure 6: Model uncertainty quantification methods on token, character and utterance levels by entropy.
@@
Equation (16) represents the uncertainty analysis at the token level. However, in speech token sequences, multiple tokens are often required to represent the pronunciation of a single character. For a character $W_{ij}$ composed of the token sequence $a_{i},a_{i+1},\dots,a_{j}$, the uncertainty $H_{\text{character}}(W_{ij})$ is computed as the average uncertainty of the tokens constituting the character. Similarly, for an utterance $S$ with a token sequence length of $|S|$, the uncertainty $H_{\text{utterance}}(S)$ is calculated as the mean uncertainty of all tokens in the utterance. The formulas for character-level and utterance-level uncertainty are as follows:
$H_{\text{character}}(W_{ij})$ $=\frac{1}{j-i}\sum_{k=i}^{j}H_{\text{token}}(P_{a_{k}})$ (17)
@@
An illustration of our detection method is provided in Figure 6. In the context of LLM generation tasks, entropy can be interpreted as a measure of the model’s uncertainty regarding its output. When the model is highly confident in predicting a specific token, the resulting probability distribution exhibits low entropy, indicating a stable generation process. Conversely, a high-entropy distribution suggests that the model is uncertain among multiple candidate tokens, leading to increased sampling variability and potential susceptibility to randomness-induced fluctuations.
@@ A.3 Detailed Fine-grained Hallucination Phenomena Analysis with Uncertainty
Based on our experimental results at the utterance level, we observe that using CER and entropy as evaluation proxies—along with external factors such as ASR errors—inevitably introduces noise. This helps explain why the obtained correlation coefficients (0.636 and 0.649) are only moderately strong. To enable a more fine-grained analysis, we further examine token-level uncertainty during the computation of utterance-level uncertainty. We also visualize token-wise uncertainty using line plots for each test utterance, allowing for detailed inspection of local variations. By leveraging temporal alignment from ASR outputs, we identified token subsequences corresponding to individual characters. For character-level uncertainty, we computed the average uncertainty of constituent tokens and generated character-wise uncertainty line plots.
@@
Figure 7: One classic case of Mispronounced/Incorrect characters, with red dashed boxes highlighting error regions.
Mispronounced/Incorrect characters In these cases, the model generates speech containing lexically incorrect terms or phonetic errors. Since each character in speech token sequences is represented by multiple tokens, such errors manifest as faulty subsequence generation, indicating hallucinatory behavior. As is visualized in Figure 7, both token-level and character-level uncertainties are significantly elevated in these cases.
@@
Repetition/Deletion Errors These kind of errors occur when the model fails to handle repetitive text, resulting in missing or extra character repetitions. As is visualized in Figure 8, some errors arise from high uncertainty (e.g., skipped characters in complex repetitions), while others stem from overconfidence—where uncertainty decreases despite errors.
Figure 9: Two classic cases of Mid-Generation Collapse, with red dashed boxes highlighting error regions. (a) A mid-generation collapse occurred in the speech token sequence (b) Starting from a certain step in the middle, the generated content becomes entirely incoherent until termination.
@@
Figure 10: One classic case of Content Divergence, with red dashed boxes highlighting error regions. The first error region includes character mispronunciations, and the second large region is inconsistent with target text although preserving semantically related contents.
Content Divergence This involves semantically meaningful but textually mismatched content, typically in long sentences.As is visualized in Figure 10, while the generated speech remains intelligible and semantically related to targets, textual alignment breaks down, which represents the occurrence of hallucinations. Relatively large uncertainty is observed despite the superficial coherence of outputs.
Figure 11: Three classic cases of Prosody errors, with red dashed boxes highlighting error regions. (a) The error region contains two characters with a gap that should be positioned between them. (b) In this error region, the separated characters should be coherently pronounced, followed by a brief pause. (c) Both larger gaps are redundant silences and should not occur in this error region.
Prosody Errors Prosody errors include misplaced pauses or redundant silences that may impair content comprehension. As is visualized in Figure 11, unlike other categories, no consistent uncertainty pattern emerges for prosodic hallucinations. Only part of high uncertainty tokens are observed in some pauses or characters nearby. This complexity suggests the need for specialized metrics tailored to prosodic representation learning.
@@ Appendix C Datasets
Evaluation Dataset Following the methodology of CosyVoice 2, we used the official seed-tts-eval test setAnastassiou et al. (2024), which contains three sub datasets, including approximately 1,000 English test samples (test-en) from the Common Voice dataset, and 2,000 Chinese test samples (test-zh) from the DiDiSpeech-2 dataset. Additionally, there are around 400 challenging samples in the case test-hard subset, which includes difficult synthesis targets such as text repetition, tongue twisters, phonetically similar texts, and commonly mispronounced characters. We use the same testing methodology and dataset as CosyVoice 2 to demonstrate the improvements made by our approach in enhancing model performance.
@@ Appendix D Implementation Details
In this experiment, we provide two sets of hyperparameter configurations, which differ primarily in the total number of steps in the Learning Rate Optimization (LRO) process. Specifically, for all experiments, the total training rounds are fixed at 15, corresponding to approximately 3500 global steps. The entire training process takes about 7 hours on four NVIDIA H100 Hopper GPUs, with each GPU utilizing around 70 GB of memory. As described in Section 3.3, to balance model performance and training stability while suppressing reward hacking, we set the minimum reward temperature to 0.825 and the maximum learning rate to 1e-5. Two configurations with slight differences in LRO are provided, where the warm-up phase is fixed at 20 steps. We adjust the remaining cosine annealing steps to achieve total LRO steps of 1500 and 2500, corresponding to more stable training and stronger model performance respectively.
It is worth noting that, given the properties of GOAT, we speculate that it should still perform effectively even with weaker or even no prompts. Reducing the prompt length or the target synthesis text length could significantly reduce GPU memory usage. We leave the exploration of more efficient training strategies for GOAT to future work.
During the inference stage, we employ the Repetition Aware Sampling (RAS) method using the default configuration from CosyVoice 2. Specifically, the top-p value is set to 0.8, top-k is set to 25, the repetition penalty window size (win_size) is 10, and the repetition penalty coefficient (tau_r) is 0.1.
@@ Appendix G Inference Latency Analysis
We use Real-Time Factor (RTF) for inference latency analysis. RTF is a widely used metric in the speech generation field to measure the efficiency of a model in terms of its processing speed. It is defined as the ratio between the time taken for the model to generate speech and the duration of the generated audio. Specifically, the RTF is calculated as: